### v1

@@ 4.5 Training Planning Policy
The planning policy determines time allocations and smoothness weights from a sequence of remaining waypoints and a current trajectory state. To handle the variable sequence length, we use the sequence-to-sequence language model from Cho et al. (2014) as a planning policy, enhanced with context-based attention as presented in Bahdanau et al. (2015). To prevent memorization, the variational autoencoder (VAE) connects the hidden state between the encoder and decoder, densifying the hidden state into the low-dimensional feature vector as described in Bowman et al. (2015). Figure 1(b) demonstrates the planning policy, which is a sequence-to-sequence model consisting of an encoder (bi-directional GRU), decoder (GRU), VAE, and attention modules. The waypoint sequence is transformed into a fixed-dimensional feature vector through the encoder, and the initial hidden state of the encoder GRU is derived from the current trajectory state $\mathcal{D\,}p$. The decoder output is then transformed into a latent time allocation $\tilde{\mathbf{x}}_{t}$ and a smoothness weight $\tilde{\mathbf{x}}_{w}$ via dedicated fully-connected layers for each fidelity level. The latent smoothness weight undergoes normalization via a softmax function, while the latent time allocation is rescaled using the average time after an exponential function is applied: $\mathbf{x}_{w}=\text{softmax}(\tilde{\mathbf{x}}_{w}),\>\mathbf{x}_{t}=T_{avg}\exp(\tilde{\mathbf{x}}_{t})/\text{dim}(\tilde{\mathbf{x}}_{t})$, where $\text{dim}(\tilde{\mathbf{x}}_{t})$ refers to the number of the remaining waypoints. The average time $T_{avg}$ is obtained by dividing the distance between remaining waypoints by the average speed of training sequences $v_{avg}$, $T_{avg}=\sum_{i=k}^{m-1}\left\lVert\tilde{p}^{i+1}_{r}-\tilde{p}^{i}_{r}\right\rVert/v_{avg}$. To improve training stability, the decoder also outputs reconstructed waypoints and is trained to minimize waypoint reconstruction loss $\mathcal{L}_{\text{Recon}}$ which is mean square error between the input and output waypoint sequences. The outputs of the planning policy are used to generate a polynomial trajectory, using the quadratic programming in (13).
@@
The planning policy is trained to maximize the expected time reduction which is obtained from the reward estimator model. We formulate a Markov decision process using the remaining waypoint as a state variable and the unnormalized output of planning policy $\tilde{x}_{t,i}$ and $\tilde{x}_{w,i}$ as an action $a_{i}$. At the end of each episode, the expected time reduction $r(\mathbf{x})$ at each fidelity level is estimated from the reward estimator. Since the reward is zero except at the end of the episode, the discounted reward is used for the $i$-th step reward: $r_{i}=\gamma^{m-i}r(\mathbf{x})$ ($i=k+1,\dots,m$), where $\gamma\in[0,1]$ is the discount factor. At the end of each training epoch, rewards are batch-wise normalized by subtracting their mean and dividing them by the standard deviation.
The proximal policy optimization (PPO) algorithm Schulman et al. (2017) is used to update the planning policy. At each training iteration, the action $\tilde{a}_{i}\sim\pi(\cdot|a_{i})$ is sampled from the normal distribution around the output $a_{i}$. We employ PPO without a value network since training a separate value network cannot adequately account for the reward bias when estimating rewards from the continually updated feasibility boundary model. The reward is estimated based on the action from the old policy $\pi_{old}$ and the next planning policy $\pi$ is updated in the sense that maximizes the PPO clipping reward:
$r_{\text{clip},i}=\min\left(r_{\pi,i}r_{i},\text{clip}(r_{\pi,i},1-\epsilon,1+\epsilon)r_{i}\right),$
$r_{\text{entropy},i}=-\log\pi(\tilde{a}_{i}|a_{i}),$
$\mathcal{L}_{\text{Policy}}={N_{\text{batch}}}^{-1}\sum_{n=1}^{N_{\text{batch}}}\sum_{i=k+1}^{m}r_{\text{clip},i}+r_{\text{entropy},i},$ (23)
where $r_{\pi,i}=\pi(\tilde{a}_{i}|a_{i})/\pi_{old}(\tilde{a}_{i}|a_{i})$ is the probability ratio between the new policy and old policy. The PPO objective function is minimized together with the reconstruction loss and the VAE variational loss to stablize the training procedure i.e. ,
@@ 4.6 Balancing Time Reduction and Tracking Error
We leverage the property that the trajectory’s shape obtained from the quadratic programming in (13) remains invariant when both time allocation and the current trajectory state are scaled. To be specific, the polynomial $\chi(\alpha\mathbf{x},\alpha^{-1}\mathcal{D\,}p,\tilde{\mathbf{p}})$ maintains a consistent shape across all scale factors $\alpha\in\mathbb{R}_{>0}$ where $\alpha^{-1}\mathcal{D\,}p$ is defined as
@@
For $\alpha>1$, this transformation slows down the trajectory without altering its shape, enabling the reference motor commands to shift towards the hovering motor speed. Building upon this transformation, we initially scale the trajectory state by multiplying the scale factor as $\mathcal{D\,}p\rightarrow\alpha\mathcal{D\,}p$ and utilize the trained policy $\pi$ to generate feasible time allocations for the scaled trajectory state: $\mathbf{x}_{t},\mathbf{x}_{w}=\pi(\tilde{\mathbf{p}},\alpha\mathcal{D\,}p)$. The final trajectory is then generated with the policy output $[\alpha\mathbf{x}_{t},\mathbf{x}_{w}]$ from the current trajectory state $\mathcal{D\,}p$. It is noteworthy that the trajectory generated with the time allocation $\alpha\mathbf{x}_{t}$ and the trajectory state $\mathcal{D\,}p$ preserves the same shape but the speed profile is scaled with $\alpha^{-1}$ compared to the trajectory with $\mathbf{x}_{t}$ and $\alpha\mathcal{D\,}p$. Without needing additional planning policy inference, this transformation effectively balances time reduction and tracking error by scaling the speed profile while preserving the trajectory’s shape.
@@ 5 Experimental Results
The proposed algorithm is evaluated by training the planning policy and testing the output of the planning policy model with flight experiments. Three different fidelity levels of evaluations are used to create the training dataset for the reward estimator. For the first fidelity level, the reference motor commands are obtained using the idealized quadrotor dynamics-based differential flatness transform presented in Mellinger and Kumar (2011). The feasible set $\mathcal{P}_{1}(\mathcal{D\,}p)$ is defined as the set of all trajectories that start from the state $\mathcal{D\,}p$ and have reference motor commands that fall within the admissible range [0, 2200] $\mathrm{r}\mathrm{a}\mathrm{d}\mathrm{/}\mathrm{s}$. In the second and the third fidelity level, the feasible set $\mathcal{P}_{2}(\mathcal{D\,}p)$ and $\mathcal{P}_{3}(\mathcal{D\,}p)$ include all trajectories that can be flown using the flight controller Tal and Karaman (2020) with less than 20 $\mathrm{c}\mathrm{m}$ position and 15 $\mathrm{d}\mathrm{e}\mathrm{g}$ yaw tracking error. The flight experiment at the second fidelity level is performed using a 6DOF flight dynamics and inertial measurement simulation Guerra et al. (2019), whereas the flight experiment at the third fidelity level is conducted using the actual vehicle. Since the planning policy generates a trajectory starting from a midpoint in a waypoint sequence, the tracking error is estimated by simulating or flying the original trajectory to the midpoint and then switching to the updated trajectory. Initially, we conduct a comprehensive analysis of the proposed algorithm in a simulated environment, utilizing the first two fidelity levels. Then, we incorporate the third fidelity level to apply the algorithm to a real-world environment, effectively refining and adapting the policy for flight experiments.
@@ 5.1 Implementation Details
[…]
@@
The training dataset is made up of $10^{5}$ sequences of five to fourteen waypoints scaled to the size of the test room ($L_{\text{space}}=[9$\mathrm{m}$,9$\mathrm{m}$,3$\mathrm{m}$]$). The parameters for minimum snap optimization in (1) are determined through grid search as follows: $\mu_{r}=1$, $\mu_{\psi}=1$, and $\rho=0$, which are identified to generate the fastest trajectories on average within our simulated evaluation setup. The evaluation time varies across fidelity levels, resulting in differing amounts of labeled data for each level. For example, evaluations at the first fidelity level take about 10 milliseconds, while the second level takes roughly 2 seconds, and the third level around 2 minutes, with times changing based on trajectory lengths. The multi-fidelity datasets are created by labeling the subset of the lower-fidelity dataset with the corresponding evaluation method, which have the following sizes: $|\mathcal{D}_{1}|=1\times 10^{5},\,|\mathcal{D}_{2}|=2\times 10^{3},\,|\mathcal{D}_{3}|=50$. All training datasets are augmented eightfold by rotating and flipping the sequences. To determine the trajectory time at each fidelity level using the line search procedure described in (5), a binary search method is employed, utilizing 10 evaluations for each sequence. During a single epoch, the planning policy is updated with 800 batches, consisting of 200 waypoint sequences. Within each batch, we evaluate 16 samples by checking reference motor commands and 1 sample using simulation. In the real-world scenarios, we additionally select 100 sequences every 100 epochs and label the output with the real-world experiments. Training is conducted in both simulated and real-world scenarios until convergence, which takes 1,800 epoch in both cases; 1,800 trajectories are thus evaluated with the real-world flight experiments.
For both simulated and real-world scenarios, policy training is conducted on an Nvidia RTX 6000 ADA GPU. To enhance the efficiency of the training process, data points designated for simulated evaluations are directed to a separate process on an external CPU server, consisting of 28 Intel Xeon CPU nodes with 72 cores, which are evaluated in parallel with the policy update procedure. Simulated evaluations account for the majority of the training time, with each epoch requiring approximately 28 hours. In real-world scenarios, flight experiments are carried out every 100 epochs, evaluating 100 trajectories over 4 hours. Across both scenarios, the total training procedure is completed over 3 weeks. Table 1 highlights the differences in computation time across fidelity levels, along with the total number of evaluations conducted at each level during the dataset generation and the reinforcement learning process.
Table 1: Evaluation time of each fidelity level and the number of evalution used during dataset generation (Dataset) and reinforcement learning (RL). $\mathcal{P}_{1}$, $\mathcal{P}_{2}$, and $\mathcal{P}_{3}$ denote the each fidelity level which use the ideal dynamics model, simulations, and real-world flight experiments, respectively. Evaluation time denote the time to evaluate a single trajectory.
@@
Computation time 75.35 $\mathrm{m}\mathrm{s}$ 2.30 $\mathrm{s}$ $\sim$ 2 $\mathrm{m}\mathrm{i}\mathrm{n}$
@@
Reinforcement learning 23,040,000 1,440,000 1,800
5.2 Evaluation in Simulated Environment
Table 2: Comparison of tracking error and trajectory time between the minimum-snap trajectories (MinSnap) and trajectories generated by the trained policy (MFRL). The policy is trained with the first and second fidelity levels in the space size $L_{\text{space}}=[9$\mathrm{m}$,9$\mathrm{m}$,3$\mathrm{m}$]$. Waypoints undergo random shifts within the range of 0 to 3 meters (Deviation - Pos) and rotations of 0 to 45 degrees (Deviation - Yaw). Pos error and Yaw error denote position and yaw tracking errors, respectively. $\Delta$ Time represents the trajectory time reduction achieved with the trained policy model relative to the minimum-snap trajectories. A positive time reduction indicates that the trained policy results in less trajectory time; for instance, MFRL produces trajectories that are 4.70% faster than the baseline when waypoints are unchanged.
@@ 5.2 Evaluation in Simulated Environment
[…]

### v2

@@ 4.5 Training Planning Policy
The planning policy determines time allocations and smoothness weights from a sequence of remaining waypoints and a current trajectory state. Figure 8 illustrates the planning policy, a sequence-to-sequence model comprising an encoder (bi-directional GRU), decoder (GRU), VAE, and attention modules. To handle variable sequence lengths, we use the sequence-to-sequence language model from Cho et al. (2014) as the main module. To prevent memorization, a variational autoencoder (VAE) connects the encoder and decoder hidden states, densifying the hidden state into a low-dimensional feature vector as described in Bowman et al. (2015). We employ content-based attention Bahdanau et al. (2015) to capture the global shape of waypoint sequences, calculating similarity between encoder and decoder hidden states, $\mathbf{h}^{\text{enc}}_{j}$ and $\mathbf{h}^{\text{dec}}_{i}$, as follows:
$\mathbf{h}^{\text{att}}_{i,j}=\text{softmax}_{j}(\mathbf{v}_{\text{att}}\tanh(\mathbf{W}_{\text{att}}[\mathbf{h}^{\text{enc}}_{j};\mathbf{h}^{\text{dec}}_{i}])),$
$\mathbf{h}^{\text{att}}_{i}=\textstyle\sum_{j}\mathbf{h}^{\text{att}}_{i,j}\mathbf{h}^{\text{enc}}_{j},$ (25)
where $\mathbf{v}_{\text{att}}$ and $\mathbf{W}_{\text{att}}$ are the weight matrix of the attention module. The non-linear activation function is chosen to optimize training stability. The waypoint sequence is transformed into a fixed-dimensional feature vector through the encoder, and the initial hidden state of the encoder GRU is derived from the current trajectory state $\mathcal{D\,}p$. The decoder output is then transformed into a latent time allocation $\tilde{\mathbf{x}}_{t}$ and a smoothness weight $\tilde{\mathbf{x}}_{w}$ via fully-connected layers (MLP). We manage separate policy models, $\pi_{l}$, for different fidelity levels, but each model shares most parameters except for the last MLP layer. To elaborate, the full architecture features one shared sequence-to-sequence backbone and three MLP heads, and this entire system—backbone and all heads—is jointly trained with the reward estimator. The latent smoothness weight undergoes normalization via a softmax function, while the latent time allocation is rescaled using the average time after an exponential function is applied:
$\mathbf{x}_{t}=T_{avg}\exp(\tilde{\mathbf{x}}_{t})/\text{dim}(\tilde{\mathbf{x}}_{t}),$ (26)
$\mathbf{x}_{w}=\text{softmax}(\tilde{\mathbf{x}}_{w}),$ (27)
where $\text{dim}(\tilde{\mathbf{x}}_{t})$ refers to the number of the remaining waypoints. The average time $T_{avg}$ is obtained by dividing the distance between remaining waypoints by the average speed of training sequences $v_{avg}$, $T_{avg}=\sum_{i=k}^{m-1}\left\lVert\tilde{p}^{i+1}_{r}-\tilde{p}^{i}_{r}\right\rVert/v_{avg}$. To improve training stability, the decoder also outputs reconstructed waypoints and is trained to minimize waypoint reconstruction loss: $\mathcal{L}_{\text{Recon}}=\lVert\tilde{\mathbf{p}}^{\text{new}}-\tilde{\mathbf{p}}^{\text{out}}\rVert^{2}$, which is the mean square error between the input and output waypoint sequences. The outputs of the planning policy are used to generate a polynomial trajectory, using the quadratic programming in (13).
@@
Figure 10: Structure of RL episode. The policy outputs time allocation and smoothness weight between consecutive waypoints. The agent receives a reward at episode end, based on estimated time reduction from all actions. The $l$-th fidelity policy uses the corresponding fidelity reward estimator.
The planning policy is trained to maximize the expected time reduction which is obtained from the reward estimator model. As illustrated in Figure, 10, we formulate a Markov decision process using the remaining waypoint as a state variable, $s_{i}$ and the unnormalized output of planning policy $\tilde{x}_{t,i}$ and $\tilde{x}_{w,i}$ as an action $a_{i}$:
$s_{i}\coloneq\tilde{p}^{\text{new}}_{i},\;\;a_{i}\coloneq[\tilde{x}_{t,i},\tilde{x}_{w,i}]$ (29)
At the end of each episode, the actual time allocation is calculated using (26), which is used to estimate the expected time reduction $r_{l}(\mathbf{x})$ from the reward estimator. $r_{l}(\mathbf{x})$ is estimated using the corresponding level’s feasibility prediction and time reduction based on its minimum snap time allocation. Since the reward is zero except at the end of the episode, the discounted reward is used for the $i$-th step reward: $r_{l,i}=\gamma^{m-i}r_{l}(\mathbf{x})$ ($i=k+1,\dots,m$), where $\gamma\in[0,1]$ is the discount factor. At the end of each training epoch, rewards are batch-wise normalized by subtracting their mean and dividing them by the standard deviation. It is noteworthy that all time allocations and smoothness weights can be considered as a single action, formulating the optimization as an RL problem with a single rollout. Instead, we frame it as a multiple rollout RL problem, which improves training stability although this formulation doesn’t strictly satisfy the Markov condition.
The proximal policy optimization (PPO) algorithm Schulman et al. (2017) is used to update the planning policy. At each training iteration, the action $\tilde{a}_{i}\sim\pi_{l}(\cdot|a_{i})$ is sampled from the normal distribution around the policy output $a_{i}$. We employ PPO without a value network since training a separate value network cannot adequately account for the reward bias when estimating rewards from the continually updated feasibility boundary model. The reward is estimated based on the action from the old policy $\pi^{\text{old}}_{l}$, and the next planning policy $\pi_{l}$ is updated in the sense that maximizes the PPO clipping reward:
$\mathcal{L}^{\text{Clip}}_{l,i}=\min\left(\omega_{l,i}\>r_{l,i},\;\text{clip}(\omega_{l,i},1-\epsilon,1+\epsilon)r_{l,i}\right),$
$\mathcal{L}^{\text{Entropy}}_{l,i}=-\log\pi_{l}(\tilde{a}_{i}|a_{i}),$
$\mathcal{L}_{\text{Policy}}=\sum\nolimits_{l=1}^{L}\sum\nolimits_{i=k+1}^{m}\mathcal{L}^{\text{Clip}}_{l,i}+\mathcal{L}^{\text{Entropy}}_{l,i},$ (30)
where $\omega_{l,i}=\pi_{l}(\tilde{a}_{i}|a_{i})/\pi^{\text{old}}_{l}(\tilde{a}_{i}|a_{i})$ is the probability ratio between the new policy and old policy. The PPO objective function is minimized together with the reconstruction loss and the VAE variational loss to stablize the training procedure i.e. ,
@@ 4.6 Balancing Time Reduction and Tracking Error
$\alpha^{-1}\mathcal{D\,}p\coloneqq\begin{bmatrix}\alpha^{-1}p_{r}^{(1)}\kern 2.0pt\cdots\kern 2.0pt\alpha^{-4}p_{r}^{(4)}\kern 2.0pt\alpha^{-1}p_{\psi}^{(1)}\kern 2.0pt\alpha^{-2}p_{\psi}^{(2)}\end{bmatrix}$. (32)
For $\alpha>1$, this transformation slows down the trajectory without altering its shape, enabling the reference motor commands to shift towards the hovering motor speed. Building upon this transformation, we initially scale the trajectory state by multiplying the scale factor as $\mathcal{D\,}p\rightarrow\alpha\mathcal{D\,}p$ and utilize the trained policy $\pi$ to generate feasible time allocations for the scaled trajectory state: $\mathbf{x}_{t},\mathbf{x}_{w}=\pi(\tilde{\mathbf{p}},\alpha\mathcal{D\,}p)$. The final trajectory is then generated with the policy output $\alpha\mathbf{x}_{t},\mathbf{x}_{w}]$ from the current trajectory state $\mathcal{D\,}p$. It is noteworthy that the trajectory generated with the time allocation $\alpha\mathbf{x}_{t}$ and the trajectory state $\mathcal{D\,}p$ preserves the same shape but the speed profile is scaled with $\alpha^{-1}$ compared to the trajectory with $\mathbf{x}_{t}$ and $\alpha\mathcal{D\,}p$. Without needing additional planning policy inference, this transformation effectively balances time reduction and tracking error by scaling the speed profile while preserving the trajectory’s shape.
@@ 5 Experimental Results
The proposed algorithm is evaluated by training the planning policy and testing the output of the planning policy model with flight experiments. Three different fidelity levels of evaluations are used to create the training dataset for the reward estimator. For the first fidelity level, the reference motor commands are obtained using the idealized quadrotor dynamics-based differential flatness transform presented in Mellinger and Kumar (2011). The feasible set $\mathcal{P}_{1}(\mathcal{D\,}p)$ is defined as the set of all trajectories that start from the state $\mathcal{D\,}p$ and have reference motor commands that fall within the admissible range, $\underline{u},\bar{u}]=[0,2200]\>$\mathrm{r}\mathrm{a}\mathrm{d}\mathrm{/}\mathrm{s}$\$. In the second and third fidelity level, the feasible sets $\mathcal{P}_{2}(\mathcal{D\,}p)$ and $\mathcal{P}_{3}(\mathcal{D\,}p)$ include all trajectories that can be flown using the flight controller Tal and Karaman (2020) with position tracking errors less than the maximum allowable value of $\bar{r}=20\>$\mathrm{c}\mathrm{m}$\$, and yaw tracking errors less than $\bar{\psi}=15\>$\mathrm{d}\mathrm{e}\mathrm{g}$\$. The flight experiment at the second fidelity level is performed using a 6DOF flight dynamics and inertial measurement simulation Guerra et al. (2019), whereas the flight experiment at the third fidelity level is conducted using the actual vehicle. Since the planning policy generates a trajectory starting from a midpoint in a waypoint sequence, the tracking error is estimated by simulating or flying the original trajectory to the midpoint and then switching to the updated trajectory. Initially, we conduct a comprehensive analysis of the proposed algorithm in a simulated environment, utilizing the first two fidelity levels. Then, we incorporate the third fidelity level to apply the algorithm to a real-world environment, effectively refining and adapting the policy for flight experiments.
@@ 5.1 Implementation Details
In the reward estimator, the input includes time allocation and smoothness weights in addition to the waypoint sequences: $x^{\text{in}}_{\text{reward},i}=[\tilde{p}^{i}_{r}/(L_{\text{space}}/2)\;\;\cos(\tilde{p}^{i}_{\psi})\;\;\sin(\tilde{p}^{i}_{\psi})\;\;f_{\text{EOS}}\;\;x_{t,i}/T_{avg}\;\;x_{w,i}]$. Similar to the planning policy, the input data is embedded into a 64-dimensional vector, with initialization of hidden states from the normalized current state and the use of identical MLP sizes. The final hidden state of the GRU is converted into a 64-dimensional feature vector via a 2-layer MLP. The MFGPC module predicts feasibility from this feature vector, Gaussian process is approximated with 128 inducing points. The regularization coefficient between fidelity levels, denoted as $c_{\text{reg}}$ in (11), is set to 1e-4. Across the encoder, decoder, and all MLPs, the hidden layer dimensions are consistently set to 256. The reward parameters are configured as follows: $r_{max}=0.1$, $r_{bias}=0.1$, the reward decay factor is $\gamma=0.9$, and the PPO clipping parameter is $\epsilon=0.2$. Both policy and reward estimator updates utilize the Adam optimizer with a learning rate of $1\text{\times}{10}^{-4}$.
The training dataset $\mathcal{D}_{init}$ is made up of $10^{5}$ sequences of five to fourteen waypoints scaled to the size of the test rooms ($L_{\text{space}}=[9$\mathrm{m}$,9$\mathrm{m}$,3$\mathrm{m}$],\>\text{and}\;[20$\mathrm{m}$,20$\mathrm{m}$,4$\mathrm{m}$\$]). The parameters for minimum snap optimization in (1) are determined through grid search as follows: $\mu_{r}=1$, $\mu_{\psi}=1$, and $\rho=0$, which are identified to generate the fastest trajectories on average within our simulated evaluation setup. Table 1 demonstrates the characteristics of waypoint sequences in the generated dataset.
Table 1: Average distance and curvature between consecutive waypoints in the training dataset, with pairwise traversal times obtained by the snap minimization method.
$L_{\text{space}}$ Distance Curvature Time
$9$\mathrm{m}$,9$\mathrm{m}$,3$\mathrm{m}$\$] 4.46 $\mathrm{m}$ 0.25 $\mathrm{m}$^{-1}$ 1.31 $\mathrm{s}$
$20$\mathrm{m}$,20$\mathrm{m}$,4$\mathrm{m}$\$] 9.68 $\mathrm{m}$ 0.13 $\mathrm{m}$^{-1}$ 1.61 $\mathrm{s}$
@@
The evaluation time varies across fidelity levels, resulting in differing amounts of labeled data for each level. Table 2 highlights the differences in evaluation time across fidelity levels, along with the total number of evaluations conducted at each level during dataset generation and MFRL training. The multi-fidelity datasets are created by labeling subsets of the lower-fidelity dataset with the corresponding evaluation method. All training datasets are augmented eightfold by rotating and flipping the sequences. To determine the trajectory time at each fidelity level using the line search procedure described in (5), a binary search method is employed, utilizing 10 evaluations for each sequence. During a single epoch, the planning policy is updated with 800 batches, consisting of 200 waypoint sequences. Within each batch, we evaluate 16 samples by checking reference motor commands and 1 sample using simulation. In the real-world scenarios, we additionally select 100 sequences every 100 epochs and label the output with the real-world experiments. Training is conducted in both simulated and real-world scenarios until convergence, which takes 1,800 epoch in both cases; 1,800 trajectories are thus evaluated with the real-world flight experiments.
Table 2: Comparison of evaluation times and evaluation counts across fidelity levels. $\mathcal{P}_{1}$ (ideal dynamics), $\mathcal{P}_{2}$ (simulation), and $\mathcal{P}_{3}$ (real-world experiments) represent each fidelity level. Evaluation time is the average per-trajectory evaluation time. Dataset size indicates labeled training sequences. Per-epoch BO subset shows the number of samples selected for reward estimator optimization. Number of evaluations represents the total evaluations during dataset generation and MFRL training.
@@
Evaluation time 75.35 $\mathrm{m}\mathrm{s}$ 2.30 $\mathrm{s}$ $\sim$ 2 $\mathrm{m}\mathrm{i}\mathrm{n}$
Dataset size $|\mathcal{D}_{1}|$ = 100,000 $|\mathcal{D}_{2}|$ = 2,000 $|\mathcal{D}_{3}|$ = 50
Per-epoch BO subset 12,800 200 1
@@
MFRL training 23,040,000 1,440,000 1,800
To enhance training efficiency, evaluations for the first and second fidelity levels are processed on an external CPU server with 28 Intel Xeon nodes and 72 cores, parallel to policy updates on an Nvidia RTX 6000 ADA GPU. Multiple trajectory evaluations are executed concurrently across available cores. The second fidelity level, including full controllers and dynamics simulation, demands significantly more memory compared to the first fidelity level, which limits the number of concurrent evaluations possible. The majority of training time, approximately 28 hours per epoch, is consumed by these evaluations, particularly those at the second fidelity level. In real-world scenarios, flight experiments are carried out every 100 epochs, evaluating 100 trajectories over 4 hours. The total training procedures across both scenarios are completed over 3 weeks, mostly due to the multi-fidelity evaluation to refine the reward estimator, while RL updates are conducted in parallel without becoming a bottleneck. Training dataset generation takes a few days as it requires a much smaller dataset, estimating the same trajectory multiple times at varying velocities.
Table 3: Comparison of tracking error and trajectory time between the minimum-snap trajectories (MinSnap) and trajectories generated by the trained policy (MFRL). The policy is trained with the first and second fidelity levels in the space size $L_{\text{space}}=[9$\mathrm{m}$,9$\mathrm{m}$,3$\mathrm{m}$\$]. Waypoints undergo random shifts within the range of 0 to 3 meters (Deviation - Pos) and rotations of 0 to 45 degrees (Deviation - Yaw). Pos error and Yaw error denote mean and standard deviation of position ($\bar{r}^{\text{err}}$) and yaw ($\bar{\psi}^{\text{err}}$) tracking errors. $\Delta$ Time represents the trajectory time reduction achieved with the trained policy model relative to the minimum-snap trajectories. A positive time reduction indicates that the trained policy results in less trajectory time; for instance, MFRL produces trajectories that are 4.70% faster than the baseline when waypoints are unchanged.
@@
0.0 $\mathrm{m}$ /   0 $\mathrm{\SIUnitSymbolDegree}$ 0.08 $\pm$ 0.02 $\mathrm{m}$ 12.8 $\pm$ 1.3 $\mathrm{\SIUnitSymbolDegree}$ 0.09 $\pm$ 0.01 $\mathrm{m}$ 12.0 $\pm$ 3.7 $\mathrm{\SIUnitSymbolDegree}$ 4.70 %
1.0 $\mathrm{m}$ / 15 $\mathrm{\SIUnitSymbolDegree}$ 0.16 $\pm$ 2.08 $\mathrm{m}$ 21.5 $\pm$ 29.4 $\mathrm{\SIUnitSymbolDegree}$ 0.09 $\pm$ 0.03 $\mathrm{m}$ 13.5 $\pm$ 10.0 $\mathrm{\SIUnitSymbolDegree}$ 4.50 %
2.0 $\mathrm{m}$ / 30 $\mathrm{\SIUnitSymbolDegree}$ 0.26 $\pm$ 2.91 $\mathrm{m}$ 40.1 $\pm$ 52.7 $\mathrm{\SIUnitSymbolDegree}$ 0.11 $\pm$ 0.07 $\mathrm{m}$ 19.2 $\pm$ 25.1 $\mathrm{\SIUnitSymbolDegree}$ 3.51 %
3.0 $\mathrm{m}$ / 45 $\mathrm{\SIUnitSymbolDegree}$ 0.36 $\pm$ 3.09 $\mathrm{m}$ 60.7 $\pm$ 66.7 $\mathrm{\SIUnitSymbolDegree}$ 0.13 $\pm$ 0.14 $\mathrm{m}$ 27.3 $\pm$ 39.2 $\mathrm{\SIUnitSymbolDegree}$ 2.49 %
5.2 Evaluation in Simulated Environment
(a) Initial and updated waypoints
(b) Reference motor speeds over waypoints
(c) Simulated motor speeds over flight time
(d) Tracking error using the initial time allocation
(e) Tracking error using the adapted time allocation
(f) Comparison of the initial speed profile
(g) Comparison of the adapted speed profile